### v1

# Understanding and Mitigating the Threat of Vec2Text to Dense Retrieval Systems

Abstract
The introduction of Vec2Text, a technique for inverting text embeddings, has raised serious privacy concerns within dense retrieval systems utilizing text embeddings, including those provided by OpenAI and Cohere. This threat comes from the ability for a malicious attacker with access to text embeddings to reconstruct the original text.
In this paper, we investigate various aspects of embedding models that could influence the recoverability of text using Vec2Text. Our exploration involves factors such as distance metrics, pooling functions, bottleneck pre-training, training with noise addition, embedding quantization, and embedding dimensions – aspects not previously addressed in the original Vec2Text paper. Through a thorough analysis of these factors, our aim is to gain a deeper understanding of the critical elements impacting the trade-offs between text recoverability and retrieval effectiveness in dense retrieval systems. This analysis provides valuable insights for practitioners involved in designing privacy-aware dense retrieval systems. Additionally, we propose a straightforward fix for embedding transformation that ensures equal ranking effectiveness while mitigating the risk of text recoverability.
Furthermore, we extend the application of Vec2Text to the separate task of corpus poisoning, where, theoretically, Vec2Text presents a more potent threat compared to previous attack methods. Notably, Vec2Text does not require access to the dense retriever’s model parameters and can efficiently generate numerous adversarial passages.
In summary, this study highlights the potential threat posed by Vec2Text to existing dense retrieval systems, while also presenting effective methods to patch and strengthen such systems against such risks.
1 Introduction
Text embeddings are dense vector representations which capture semantic information about the text they encode Muennighoff et al. (2023). Search engines that leverage these embeddings often employ dense retrievers (DRs) Tonellotto (2022); Zhao et al. (2022); Guo et al. (2022); Bruch (2024). These retrievers utilize text embedding models to encode both queries and documents into embeddings. A similarity metric, such as cosine similarity, is then used to estimate relevance. DRs have demonstrated improved retrieval effectiveness compared to traditional exact term-matching search systems, arguably due to the rich semantic information encoded in the embeddings Yates et al. (2021).
However, a recent study conducted by Morris et al. (2023) raises serious privacy concerns regarding DRs. This study explored the issue of inverting textual embeddings: recovering the original text from its embedding. The proposed Vec2Text method iteratively corrects and generates text to reconstruct the original text based on the given input embedding. Vec2Text can accurately recover 92% of short text and reveal sensitive information (such as patient names in clinical notes) with high accuracy.
Even more concerning is that training Vec2Text does not require access to the embedding model parameters; only the text-embedding pairs from the training data are needed. This implies that DR systems that use popular text embedding API services, such as those provided by OpenAI and Cohere, can be vulnerable to attacks from malicious attackers. These attackers can build their own training data by sending numerous text snippets to the API and obtaining the corresponding embeddings. Subsequently, these text-embedding pairs can be utilized to train a Vec2Text model against the target embedding model. Once the Vec2Text model is trained, the document embeddings stored in these DR systems are no longer secure. An attacker with access to the document embeddings can recover the original text from the embeddings using the trained Vec2Text model.
A limitation of the original Vec2Text paper was that it investigated only two embedding models [^1] GTR-base and OpenAI text-embeddings-ada-002.. Presently, there are numerous DR systems that utilize various types of embedding models. The distinctions among these embedding models arise from different strategies employed during training or inference, such as distance metrics, pooling functions, bottleneck pre-training, embedding quantization, and embedding dimensions. These diverse strategies involve trade-offs between retrieval effectiveness and efficiency for DRs. However, the impact of these strategies on Vec2Text reconstruction remains unexplored. In this paper, we address this gap by re-training Vec2Text models against various DR systems that implement these different strategies. Our comprehensive experimental results reveal that some strategies can make DRs more susceptible to Vec2Text, while others effectively protect privacy without compromising retrieval effectiveness. Additionally, we propose a simple yet effective embedding transformation method, secret to each user, which ensures that retrieval effectiveness remains intact while completely mitigating the risk of text reconstruction. Our study provides valuable insights for practitioners designing privacy-aware dense retrieval systems.
Our in-depth look at Vec2Text further revealed another vulnerability affecting dense retrievers: Vec2Text can be employed to conduct corpus poisoning attacks on dense retrievers Zhong et al. (2023). A corpus poisoning attack involves a malicious actor generating adversarial passages to maximize similarity with a set of training queries. Adversarial passages are then inserted into a retrieval corpus, effectively tricking the ranker into retrieving them for all unseen user queries, thus undermining the user experience of the targeted search system. Vec2Text presents a theoretically more dangerous method for corpus poisoning compared to previous approaches because it does not require access to the embedding model parameters and can easily generate a large number of adversarial passages. However, there has been no study yet to investigate how Vec2Text performs in corpus poisoning attacks. Therefore, in this paper, we present our preliminary results of applying Vec2Text to the corpus poisoning task. Our findings demonstrate that Vec2Text could pose a serious threat to current DR systems.
Our contributions can be summarized in five main aspects:
  • We reproduce Vec2Text, fixing a serious experimental error in the use of the output layer of the GTR-base embeddings. This fix has implications in terms of both retrieval and reconstructibility effectiveness, although the trends observed in the original work do not change.
  • We re-train Vec2Text against a wide range of embedding models to understand the key elements that contribute to the text reconstruction.
  • Our comprehensive results provide insights into the trade-offs between privacy protection and retrieval effectiveness.
  • We propose an embedding transformation method to mitigate privacy risks without compromising retrieval performance.
  • We are the first to adapt Vec2Text to the corpus poisoning task, empirically demonstrating Vec2Text’s capability to efficiently generate adversarial passages without access to embedding model parameters, posing a serious threat to DR systems’ integrity and user experience.
2 The Vec2Text Method
Figure 1: Overview of Vec2Text, taken from Morris et al. (2023) with permission. “Given access to a target embedding $e$ (blue) and query access to an embedding model $\phi$ (blue model), the system aims to iteratively generate (yellow model) hypotheses $\hat{e}$ (pink) to reach the target. Example input is a [passage] taken from a recent Wikipedia article (June 2023). Vec2Text perfectly recovers this text from its embedding after 4 rounds of correction.”
Vec2Text inverts the text encoding process: given an input embedding, generate the text it represents Morris et al. (2023). (A graphical overview of Vec2Text, taken from Morris et al. (2023) with permission, is shown in Figure 1.) The method comprises two training stages. In the initial stage, a hypothesis text generation model is trained, utilizing a conditional transformer generative model that exclusively takes the embedding as the model input. The training objective is to produce the original text. However, Morris et al. found that this simplistic model is insufficient for generating highly accurate original text. Consequently, they consider the text generated by this first stage model as a hypothesis.
Moving on to the second stage, the learning process involves training another transformer generative model. This model aims to generate satisfactory text by refining the initial hypothesis generated in the first stage. The refinement is achieved through an iterative re-embedding and correction training objective. In each iteration step, the model takes the ground-truth embedding, the generated text, and its embedding from the last iteration step as inputs (the 0 step uses the hypothesis text generated from the first stage model). The output target is the original text. This iterative process allows the model to focus on the differences between the generated text and the original text in the embedding space and gradually reduce these differences.
The models in Vec2Text are parameterized as a standard encoder-decoder transformer conditioned on the previous output. One challenge is inputting conditioning embeddings into the transformer encoder, which requires a sequence of embeddings as input. To address this, a small multi layer perceptron is used to project a single embedding vector to a larger size and reshape it to match the input requirements of the encoder.
During inference, beam search is used to guide the generation. At each step of correction, a number of possible corrections are considered, and the top unique continuations are selected based on their distance in embedding space to the ground-truth embedding.
Morris et al. (2023) trained Vec2Text to invert222i.e. to reconstruct the original text that corresponds to the embeddings. GTR-base Ni et al. (2022) and OpenAI’s text-embeddings-ada-002 embedding models. It achieved high reconstructibility333i.e. the capability to reconstruct the original text from the embedding. on in-domain datasets and adapted well to different-length inputs. The method outperformed baselines on various metrics, and running beam search improved the exact match score. In a case study on MIMIC III clinical notes, Vec2Text demonstrated high performance in reconstructing GTR-embedded clinical notes, re-identifying high percentages of first names, last names, and complete names: this raises serious privacy concerns about the embeddings.
However, we note that Morris et al. only considered reconstructing GTR-base and OpenAI text-embeddings-ada-002 embedding models. Many other different embedding models exist, and each uses very different training and inference strategies. Can all embedding models be easily attacked by Vec2Text? What types of embedding models are vulnerable to Vec2Text? In the following sections, we explore a wide range of key decisions in constructing embedding models (e.g., pooling functions, bottleneck pre-training, quantization), that go well beyond the scope of the original paper.
3 Reproduction of vec2text
3.1 Experimental Methodology
We begin by completing a full reproduction of Vec2Text. This is needed to ensure a fair comparison of the different embeddings models and settings with which we further study Vec2Text.
We use the official Vec2Text codebase [^4] https://github.com/jxmorris12/vec2text provided by the authors and adhere to the training parameters outlined in the original Vec2Text paper. Specifically, we train Vec2Text on documents from the Natural Question (NQ) dataset Kwiatkowski et al. (2019), setting the max sequence length to 32. The Adam optimizer is employed with a learning rate of $2\times 10^{-4}$, incorporating warmup and linear decay. A batch size of 512 is utilized, and all models are trained on a single H100 GPU. The only deviation from the original setting is the number of training epochs. We note that the complete Vec2Text training pipeline is time-consuming. The authors reported in the paper that the full pipeline takes 2 days to train for 100 epochs with four A6000 GPUs. However, in our implementation using the provided code, the full pipeline takes approximately 10 days with a single H100 GPU. Given the extensive number of embedding models we plan to investigate and the limited computational resources at our disposal, we train for 50 epochs. This reduces the training time to around 5 days for each embedding model, making the process more feasible within our constraints. For the Vec2Text inference, we always use beam search with 50 steps and a beam width 4.
In addition to the variation in the number of epochs, we have identified a significant issue in the original experiment involving the training of Vec2Text for reconstructing text from the GTR-base embeddings. Morris et al. inadvertently used the last hidden layer outputs of the GTR-base model as the final text embeddings. However, it’s crucial to note that GTR-base incorporates an additional dense pooling layer to project the last hidden layer outputs, a detail overlooked by the Vec2Text training code. To address this issue, this section includes a re-training of Vec2Text with the correct GTR-base embeddings, and we compare the results with the original setting. This correction ensures a more accurate evaluation of Vec2Text’s performance in inverting GTR-base embeddings.
Finally, to compare how different embedding models balance retrieval effectiveness and privacy preservation, we also measure the retrieval effectiveness of all considered embedding models. Following the original Vec2Text training setting, which involves training Vec2Text on documents from the NQ dataset, we adopt the top-k retrieval accuracy evaluation setting from DPR Karpukhin et al. (2020), which measures the percentages of retrieved passages that contain a span that answers the question (query). This evaluation is commonly employed in the literature for assessing retrieval effectiveness on the NQ dataset Gao and Callan (2022); Wang et al. (2023); Qu et al. (2021); Ma et al. (2021); Xiong et al. (2020). We use the official NQ test queries and Pyserini Lin et al. (2021) IR toolkit for this evaluation.
To evaluate reconstructibility, i.e. the ability to recover the original text from its embedding, we adopt the same measures used by Morris et al. (2023):
  • BLEU score Papineni et al. (2002), a metric capturing n-gram similarities between the true and reconstructed text;
  • Token F1, representing the multi-class F1 score between the predicted and true token sets;
  • Exact-match, indicating the percentage of reconstructed outputs perfectly matching the ground truth;
  • Cosine similarity, measuring the similarity between the true embedding and the embedding of the reconstructed text.
We randomly sample 1,000 passages from a held-out set of NQ training data for this evaluation [^5] From their code, we understood that also Morris et al. performed sampling; while they do not mention how many passages they sampled, the default value in their code is 1,000 and so we adopted that value..
Table 1: Reproduction results of Vec2Text Morris et al. (2023) comparing retrieval effectiveness and text reconstructibility. Bug fixed represents a correction to ensure Vec2Text uses the final dense pooling layer of GTR-base for output. Subscripts denote statistically significantly difference in retrieval effectiveness using a paired, two-tails t-test ($p<0.01$).
Embedding Models Retrieval effectiveness Reconstructibility
top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. GTR-base (original) 0.634 0.711 0.816 0.893 90.128 0.970 0.758 0.990
b. GTR-base (reproduced Morris et al. (2023)) 0.634 0.711 0.816 0.893 88.956 0.971 0.699 0.996
[…]
3.2 Reproduction Results
Table 1 shows Vec2Text reproduction results. The first row in the table shows the results obtained using the Vec2Text model checkpoint provided by the authors, which was trained with incorrect GTR-base embeddings. Despite the incorrect embeddings, the results indicate that it still performs well for retrieval, yielding reasonable top-k retrieval accuracy. Additionally, Vec2Text demonstrates robust reconstructibility. It is noteworthy that these outcomes are achieved with the incorrect GTR-base embeddings (without the dense pooling layer), suggesting that the last hidden layer of GTR-base already contains sufficient information for Vec2Text to effectively reconstruct the original input text.
The second row in the table shows the results obtained by our reproduced Vec2Text model with incorrect GTR-base embeddings (i.e., trying to exactly reproduce the original Vec2Text results). The only difference from the first row is the number of epochs; our reproduction uses half the number of epochs (50 epochs) compared to the original checkpoint (100 epochs). Since our goal is to replicate the original Vec2Text results, we also utilize the incorrect GTR-base embeddings, resulting in exactly the same retrieval effectiveness to the original setting. Regarding Vec2Text reconstructibility, our reproduced checkpoint only shows a slight decrease in the exact match score but achieves very close scores on other evaluation metrics. These results suggest that Vec2Text almost converges with half of the epochs than those used by Morris et al., and the only notable improvement with longer training is in exact reconstruction. From here we only consider Vec2Text models trained on 50 epochs, noting their results are likely similar to those obtained if 100 epochs were used.
Finally, in the last row, we present our results using the correct GTR-base embeddings for retrieval and training Vec2Text with 50 epochs [^6] We have made the checkpoint publicly available on the HuggingFace: https://huggingface.co/ielabgroup/vec2text_gtr-base-st_corrector . Our findings emphasize the significance of using the correct embedding for retrieval, as the top-k accuracy exhibits a substantial increase compared to that obtained with incorrect embeddings. On the other hand, we achieve comparable Vec2Text reconstructibility with the original checkpoint, despite utilizing only half the training epochs. These results suggest that employing the correct outputs from the embedding model, rather than the intermediate layer outputs, may contribute to improved Vec2Text reconstruction.
Table 2: Reproduction of the DPR Karpukhin et al. (2020) embedding models with different distance metrics and pooling functions. "cls" indicates the utilization of the CLS token embedding as the representation, while "mean" indicates the use of mean pooling, aggregating all token embeddings to generate the final text embedding. "dot" indicates the dot product distance metric and "cos" indicates the cosine similarity distance metric. There is no statistically significant difference observed in terms of retrieval effectiveness.
Embedding Models Retrieval effectiveness Reconstructibility
Label Metric Representation top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. DPR_cls_dot Dot-product CLS token 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_cos Cosine similarity CLS token 0.748 0.799 0.863 0.914 71.834 0.879 0.345 0.975
c. DPR_mean_cos Cosine similarity Mean pooling 0.745 0.802 0.866 0.916 81.918 0.924 0.480 0.988
4 Understanding what impacts Vec2Text effectiveness
In the previous sections, we successfully reproduced Vec2Text with the GTR-base embedding model and finalized the experimental environment. In this section, our goal is to understand how various strategies implemented by different embedding models impact retrieval effectiveness and Vec2Text reconstructibility. To achieve this, we re-train Vec2Text with respect to the target embedding models. Each of the embedding models represents different choices in terms of distance metric, pre-training strategy, etc. When re-training, we used the same Vec2Text parameters reported above.
To maintain control over experimental parameters and prevent results from being influenced by factors such as different datasets and training parameters, we independently train the embedding models ourselves. This is done with the same experimental settings but utilizing different training or inference strategies. Specifically, we start by reproducing the DPR embedding model Karpukhin et al. (2020), following the training parameters outlined in the DPR paper and employing the Tevatron dense retriever training toolkit Gao et al. (2023) to train the DPR model on the NQ dataset, the same dataset used to train Vec2Text. After reproducing the DPR retrieval results, we then apply the same dense retriever training hyperparameters (e.g., batch size, learning rate, and number of hard negatives) to train other DPR variations with distinct training strategies, as detailed in the next subsections.
4.1 Distance Metric and Pooling Method
In Table 2, we examine the impact of different distance metrics, namely dot product (dot) or cosine similarity (cos), used for training and inference in DPR models. Additionally, we investigate the pooling method employed to construct a single embedding for each input text, specifically, CLS token embedding (cls) or mean pooling of all token embeddings (mean).
The DRP_cls_dot represents our reproduction of the original DPR embedding model, utilizing CLS token embedding to represent the entire text and employing the dot product during training. We achieved very similar top-k retrieval effectiveness as reported in the original DPR paper, indicating successful reproduction of DPR. It is worth noting that our trained DRP_cls_dot exhibits higher retrieval effectiveness than GTR-base embedding models presented in Table 1, possibly due to our DPR training and testing data originating from the same data distribution, while GTR-base has training data from other domains.
However, concerning Vec2Text reconstructibility, we observe that Vec2Text trained with the DRP_cls_dot embedding model yields much lower reconstructibility compared to GTR-base embedding models. This suggests that the embeddings provided by DRP_cls_dot contain less information for Vec2Text to reconstruct the original text, indicating a potentially more secure embedding model than GTR-base.
We then proceeded to train two additional DPR variations, namely DRP_cls_cos, which employs CLS token embedding and cosine similarity, and DRP_mean_cos, which utilizes mean pooling of all token embeddings with cosine similarity. There was no statistically significant differences in retrieval effectiveness for these settings.
Regarding reconstructibility, a comparison between DRP_cls_dot and DRP_cls_cos suggests that cosine similarity appears to enhance Vec2Text’s exact matching score. However, when comparing DRP_cls_cos and DRP_mean_cos, the reconstructibility of Vec2Text significantly increases. This surprising result suggests that mean pooling is a key factor for Vec2Text reconstructibility but does not necessarily contribute to improved retrieval effectiveness. We suspect that this might be due to mean pooling explicitly gathering all the token information into the embedding, whereas the CLS token embedding only implicitly learns the text information during DR training, making it less helpful for Vec2Text training. Consequently, practitioners in the field of DR should carefully consider avoiding the use of mean pooling to enhance the security of the embedding model against Vec2Text attacks.
It is worth noting that the GTR-base model exclusively employs mean pooling and cosine similarity. This may explain why Vec2Text achieves high reconstructibility scores when trained with GTR-base embedding models.
Table 3: Zero-shot retrieval and bottlenecked pre-training results. Subscripts denote statistically significantly better retrieval effectiveness with a T-test ($p<0.01$). Results show that even zero-shot retrievers are at risk to Vec2Text reconstruction. Bottleneck pre-training proved useful at improve retrieval effectiveness but greatly increased risk of reconstructibility.
Embedding Models Retrieval effectiveness Reconstructibility
Label Setting top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. BERT_cls_dot Zero-shot 0.101 0.153 0.288 0.467 37.034 0.671 0.042 0.968
[…]
4.2 Zero-shot Regime and Bottleneck Pre-training
In the previous section, we delved into fine-tuned DPR embedding models. In this section, we further explore zero-shot dense retriever with pre-trained, zero-shot embedding models. The aim is to investigate if the DR fine-tuning (done in the previous sections) itself leads to vulnerability from Vec2Text. Additionally, we explore embedding models that leverage bottleneck pre-training designed to enhance the CLS token embedding. The results are presented in Table 3.
BERT_cls_dot and BERT_mean_cos use the same backbone pre-trained embedding model (BERT Kenton and Toutanova (2019)) and perform the exact same inference as DRP_cls_dot and DRP_mean_cos but without DPR training; they represent zero-shot retrievers. Their zero-shot retrieval effectiveness is significantly worse than fine-tuned DPR models (as might be expected for a zero-shot dense retriever). Mean pooling with cosine similarity exhibits higher zero-shot retrieval effectiveness than dot product with CLS token embeddings but still well below the trained DPR models from the previous section.
Now considering reconstructibility, BERT_cls_dot is much less vulnerable than than DRP_cls_dot. This suggests that DR fine-tuning injects information that could aid Vec2Text in reconstruction from the CLS token embedding. In contrast, BERT_mean_cos already demonstrates strong Vec2Text reconstructibility, further confirming that mean pooling with cosine similarity is potentially problematic — even for the zero-shot setting.
We then tested the SimLM Wang et al. (2023) embedding model, which leverages a bottlenecked pre-training approach Gao and Callan (2021); Lu et al. (2021); Liu and Shao (2022); Shen et al. (2022); Chuang et al. (2022); Wu et al. (2022); Zhuang et al. (2023). Bottlenecked pre-training is a task designed to enhance the CLS token embedding. This is achieved by taking the embedding model’s outputted CLS token embedding and inputting it into a weaker decoder model. The pre-training objective is to enable the decoder model to perform the MASK language modelling task or, in the case of SimLM, the ELECTRA pre-training task Clark et al. (2019). Consequently, the model must learn to inject useful information into the CLS token embedding to assist the decoder in completing the task. In fact, this pre-training approach closely resembles Vec2Text training itsself, as both involve a generator that takes an embedding as input and attempts to reconstruct a piece of text. Previous studies on bottlenecked pre-training have shown improvements in DR retrieval effectiveness. However, we suspect that this pre-training approach may make the embedding particularly vulnerable to Vec2Text, raising potential privacy concerns.
Our results confirmed our hypothesis: SimLM_cls_dot with additional bottlenecked pre-training on the NQ datasets [^7] https://huggingface.co/intfloat/simlm-base-wiki100w, demonstrated higher zero-shot retrieval effectiveness than BERT_mean_cos, indicating that bottlenecked pre-training injected useful information into the CLS token that can aid in retrieval. However, the enhanced CLS token embedding proved highly vulnerable to Vec2Text. The BLEU, TF1, and COS scores were even higher than those achieved by Vec2Text trained with correct GTR-base embeddings. We conducted further tests by using SimLM as the backbone model instead of BERT to train a DPR model with the same training parameters as DPR_cls_dot. The results are outlined in the row labeled SimLM_cls_dot_finetuned. Indeed, the use of SimLM further improved DPR retrieval effectiveness and also achieved very high reconstructibility. We conclude that practitioners should carefully consider the privacy implications of bottleneck pre-training before incorporating this approach.
Table 4: Impact of embedding dimension and quantization. Both dimension reduction and product quantization prove to be simple and effective methods to mitigate Vec2Text reconstruction, without comprising retrieval effectiveness. There is no statistically significant difference observed in terms of retrieval effectiveness.
Embedding Models Index size Retrieval effectiveness Reconstructibility
Label Dim. Quantization top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. DPR_cls_dot (768) 768 None 61GB 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_dot (256) 256 None 21GB 0.731 0.784 0.855 0.910 25.060 0.565 0.035 0.881
c. DPR_cls_dot (PQ_768) 768 Product 16GB 0.749 0.801 0.862 0.914 1.916 0.146 0.000 0.748
d. DPR_cls_dot (PQ_256) 256 Product 5GB 0.740 0.796 0.864 0.912 2.305 0.159 0.000 0.767
4.3 Embedding Dimensionality & Quantization
Next, we explore how embedding dimensionality and quantization, popular methods for reducing DR index size and improving retrieval efficiency, impact retrieval effectiveness and Vec2Text reconstructability.
Regarding embedding dimensionality, we trained another DPR model with the exact settings as DPR_cls_dot, but adding a dense pooling layer on top to reduce the embedding dimension from 768 to 256. This variant is labeled as DPR_cls_dot (256). For embedding quantization, we applied product quantization (PQ) Jegou et al. (2010) to DPR_cls_dot. We experimented with two settings: one with the number of sub-vectors set to 768 and another with 256. The number of bits per sub-vector for both settings is set to 8 in both cases, resulting in embedding dimensions of 768 and 256. We label them DPR_cls_dot (PQ_768) and DPR_cls_dot (PQ_256) respectively.
The results presented in Table 4 indicate that dimensionality reduction and product quantization can significantly reduce the dense vector index sizes, as expected. Product quantization demonstrates robust retrieval effectiveness; when the sub-vector is set to 768, there is no decrease in top-k accuracy, and only a slight decrease in retrieval effectiveness when further reducing the dimensionality to 256. These findings align with previous studies Ma et al. (2021).
On the other hand, simply reducing the embedding dimension by adding a dense pooling layer can considerably decrease Vec2Text reconstructibility, albeit at the cost of hurting retrieval effectiveness. PQ once again demonstrates superior privacy protection ability. Both settings of PQ we considered completely mitigate Vec2Text reconstructibility. Notably, this is achieved without compromising retrieval effectiveness, and the index size is significantly reduced. Hence, we conclude that product quantization is a simple and effective method for protecting embeddings from attacks by Vec2Text.
Table 5: Impact of adding noise and customized embedding transformation. Noise injection was an effective reconstruction mitigation strategy. The proposed vector transformation mitigation strategy is easily applied and guaranteed to maintain retrieval effective while completely degrading reconstructibility. There is no statistically significant difference observed in terms of retrieval effectiveness.
Embedding Models Retrieval effectiveness Reconstructibility
Label Mitigation top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. DPR_cls_dot None 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_dot_noise Noise injection 0.730 0.790 0.856 0.911 10.022 0.415 0.000 0.823
c. DPR_cls_dot_transform Vector transformation 0.748 0.800 0.863 0.914 1.178 0.065 0.000 0.443
5 Mitigation Strategies
In the previous sections, we explored how different DR training and inference strategies trade-off retrieval effectiveness and embedding privacy protection. In this section, we evaluate methods that proactively protect against the Vec2Text attacks.
5.1 Noise Injection
In the original Vec2Text paper, Morris et al. explored a basic mitigation strategy of adding Gaussian noise directly to each embedding:
$\phi_{noisy}(x)=\phi(x)+\lambda\cdot\epsilon,\epsilon\sim\mathcal{N}(0,1),$ (1)
where $\phi(x)$ is the embedding for a given input text $x$, $\epsilon$ is the noise vector whose elements are sampled from a Gaussian distribution, and $\lambda$ is a hyperparameter controlling the amount of injected noise. Morris et al. showed that setting $\lambda=0.1$ completely prevents reconstructibility for the GTR-base embedding model, but it also has a major negative impact on retrieval effectiveness.
Table 6: Results of adding noise during Text2Text training. The target embedding model is $DPR\_cls\_dot$. Noise injection proved to be an effective mitigation strategy even when Vec2Text is actually trained with noisy embeddings.
Vec2Text Models Reconstructibility
On original embeddings On noisy embeddings
Label Trained on bleu tf1 exact cos bleu tf1 exact cos
Vec2Text Original embeddings 70.735 0.875 0.319 0.985 10.022 0.415 0.000 0.823
Vec2Text_noise Noisy embeddings 29.264 0.635 0.012 0.966 17.417 0.504 0.005 0.882
Figure 2: Impact on retrieval effectiveness and reconstructibility of different amounts of noise injection with different retrieval models. Larger $\lambda$ signifies more noise injection.
Here, we reproduce the noise injection mitigation strategy on our trained $DPR\-\_cls\_dot$ embedding model, labeled as $DPR\_cls\_dot\_noise$ in Table 5. Our results show partially different outcomes: setting $\lambda=0.1$ did indeed prevent Vec2Text reconstructibility but only slightly impacted retrieval effectiveness. This may be due to the sensitivity of the specific embedding models to the amount of noise (as controlled by $\lambda$). To better understand this we completed a sweep of $\lambda=[0.001,0.01,0.1,1.0]$ for DPR, GTR and GTR with corrected embeddings; this is shown in Figure 2. We observe that DPR requires more noise to see an effect: both in terms of retrieval effectiveness deterioration, and reconstructibility robustness. We hypothesise that this is due to the different distance metric used by the two methods (GTR uses cosine, while DPR here uses dot product). Overall, adding noise remains a desirable mitigation strategy but correct parameter setting is critical. However, the correct setting of the noise parameter appears to be dependent on the specific embedding model used. In the next section we propose a new mitigation strategy that does not have this drawback.
A question still remains, though, of whether adding noise was an effective mitigation strategy if Vec2Text was actually trained using these noisy vectors. To investigate this, we re-trained a Vec2Text model while adding the same Gaussian noise to the input embeddings. The results are presented in Table 6. Our findings demonstrate that adding noise to Vec2Text training reduces reconstructibility on original embeddings. While it does make Vec2Text more robust to noisy embeddings, the level of reconstructibility remains very low. Thus, we conclude that adding noise to Vec2Text training does not make the noise-adding defence vulnerable.
5.2 Embedding Transformation
Finally, we propose a new mitigation strategy that prevents reconstruction via Vec2Text but has the benefit of theoretically guaranteeing no change in retrieval effectiveness and not requiring parameter tuning that is dependent on the embedding model used.
Our idea is straightforward: we propose applying a simple transformation to all the embeddings, with a transform function customized and only visible to each user (similar to an API key) who sends text to the embedding model APIs (such as OpenAI users sending text to OpenAI APIs to obtain embeddings). Formally, we define the transformed embeddings as:
$\phi_{\text{transformed}}(x)=f(\phi(x)),$ (2)
where $f$ is a linear transformation function unique to each user. For example, in our experiments, we simply set $f(\phi(x))=-2.6\cdot\phi(x)$ so all embeddings are adjusted by a constant value. The function (in our case the constant value) is kept secret and thus not known to a Vec2Text attacker. This means that attackers will obtain different embeddings from the API calls to train their Vec2Text model. Additionally, on the user side, since the same transformation function is applied to all embeddings retrieval will not be impacted. Users can still reverse engineer to get the original embeddings since the transformation function is visible to the user.
We present our empirical results in Table 5, labeled with DPR_cls_dot_transform (row c.). Note that the retrieval effective is identical to that of DPR_cls_dot (row a.), as guaranteed by the uniform transformation. At same time, reconstructibility is almost completely degraded so privacy is protected. We believe this simple strategy could be easily adopted by service providers like OpenAI or Cohere to offer their users the option to make their embeddings more secure.
6 A New Threat: Corpus Poisoning
In this section, we investigate whether Vec2Text poses a threat to existing DR systems through a corpus poisoning attack.
6.1 Existing Corpus Poisoning Attack Method
A corpus poisoning is an adversarial attack where a malicious user injects adversarial passages into a retrieval corpus with the intention of misleading DR models. This is a relatively new type of attack, and, to date, there is only one existing study conducted on this problem Zhong et al. (2023). Specifically, Zhong et al. (2023) proposed a gradient-based approach, inspired by the HotFlip method Ebrahimi et al. (2018); Wallace et al. (2019), which generates a small set of adversarial passages by perturbing discrete tokens in randomly initialized passages to maximize their similarity with a provided set of training queries. These adversarial passages are then inserted into the retrieval corpus, and the success of the attack is determined by the retrieval of these adversarial passages at a high rank in response to future unseen queries. These adversarial passages can be used to harm retrieval effectiveness and/or inject spam or content with misinformation into the search engine result list.
The current HotFlip-based corpus poisoning method has two drawbacks (that is, from an attacker perspective). Firstly, being a gradient-based method, it requires access to the embedding model weights, i.e., it is a white box attack. Consequently, attackers cannot employ this method to target DR systems built with closed-source embedding models, such as OpenAI embedding models. Secondly, the method is iterative, with only one token in the adversarial passage selected and perturbed at each iteration. This process cannot be parallelized within each adversarial passage, as the perturbation of the next token depends on the previously perturbed token. This makes the method time-consuming and resource intensive: in our experiments with a single H100 GPU, HotFlip takes approximately 2 hours to generate a single adversarial passage.
6.2 Performing Corpus Poisoning with Vec2Text
In contrast, Vec2Text does not suffer from the drawbacks of HotFlip. First, Vect2Text does not require access to the embedding model weights. Once Vec2Text is trained, it can directly generate the adversarial passage from the embedding using standard language model generation inference. Second, it allows for efficient generation of a large set of adversarial passages (being far less computationally intensive than HotFlip).
To formally define corpus poisoning with Vec2Text, let $\mathcal{Q}=\{q_{1},q_{2},...,q_{|\mathcal{Q}|}\}$ be a set of training queries. The goal is to use Vec2Text to generate an adversarial passage $a$ whose embedding maximizes the similarity to all the training query embeddings:
$a$ $=\underset{a^{\prime}}{\mathrm{argmax}}\frac{1}{\mathcal{|Q|}}\sum_{q_{i}\in\mathcal{Q}}\phi(q_{i})^{\intercal}\phi(a^{\prime})$ (3)
$=\underset{a^{\prime}}{\mathrm{argmax}}\>\phi(a^{\prime})^{\intercal}\frac{1}{\mathcal{|Q|}}\sum_{q_{i}\in\mathcal{Q}}\phi(q_{i})$ (4)
$=\underset{a^{\prime}}{\mathrm{argmax}}\>\phi(a^{\prime})^{\intercal}\phi_{\mathcal{Q}},$ (5)
where $\phi_{\mathcal{Q}}$ is the average embedding or centroid embedding of all the training query embeddings. The maximum similarity is achieved when $\phi(a^{\prime})=\phi_{\mathcal{Q}}$. Vec2Text can solve this optimization problem through a three-stage process:
  • The first stage is the standard training of Vec2Text (i.e., reconstructing the text for any given embedding).
  • The second stage involves computing the centroid embedding $\phi_{\mathcal{Q}}$ of training queries, which can be done by sending the training queries to the embedding model to obtain all the query embeddings and averaging them.
  • Finally, inputting $\phi_{\mathcal{Q}}$ to the trained Vec2Text yields a generated text, which is the adversarial passage $a$.
A perfect Vec2Text would generate an adversarial passage whose embedding is exactly the same as the query centroid; i.e., $\phi(a)=\phi_{\mathcal{Q}}$.
Once again, we emphasize that the entire process does not require access to model weights. Moreover, generating an adversarial passage with Vec2Text is efficient: generating a single passage took 5 seconds on a single H100 GPU, employing beam search with 50 steps and a width of 4.
Table 7: Results of applying Vec2Text to the task of corpus poisoning, compared to existing HotFlip approach. Vec2Text has the advantage of being able to produce numerous adversarial passages ($k$) without model weights. Both Vec2Text and HotFlip are well below the upper bound, indicating that better methods could pose a serious risk in corpus poisoning.
Attack method #Adversarial Passages / Success@10 Success@20 Success@100 Success@1000
Query Clusters $k$
HotFlip Zhong et al. (2023) 10 0.006 0.015 0.105 0.532
Vec2Text 10 0.006 0.012 0.036 0.101
   Query centroid (upper bound) 10 0.206 0.304 0.575 0.926
HotFlip 100 $\cdots$ Too computationally expensive $\cdots$
Vec2Text 100 0.061 0.091 0.189 0.401
   Query centroid (upper bound) 100 0.493 0.620 0.868 0.994
HotFlip 1000 $\cdots$ Too computationally expensive $\cdots$
Vec2Text 1000 0.268 0.335 0.521 0.791
   Query centroid (upper bound) 1000 0.779 0.873 0.980 1.000
6.3 Corpus Poisoning Experiments
$a$ = "That drawings promises Key kot iner jor Respond machines AK <pad> ance pe izi very nie <pad> <pad> Doar ceapa Weg MEN Am OK Jamie words Um Imp refers ’’ War Hop nism the serving Bol auto Palatul fries rome lighter (1 49 fetch 19 cities can counting 16, As"
(a) HotFlip. Cosine similarity of this passage to the query centroid embedding is $\cos(\phi(a),\phi_{\mathcal{Q}})=0.96$.
$a$ = "when does the 7 episode season of the new two come out </s>"
(b) Vect2Text. Cosine similarity of this passage to the query centroid embedding is $\cos(\phi(a),\phi_{\mathcal{Q}})=0.85$.
Figure 3: Sample adversarial passages most closely matching the NQ training query centroid.
We conducted the corpus poisoning experiments with the GTR-base embedding model (with correct embeddings) as the target DR system, utilizing the NQ datasets as the target corpus. NQ training queries are used to first encode the query centroid embedding and then to generate adversarial passages. In practice, we generate multiple adversarial passages (in line with Zhong et al. (2023)). This was done by first using $k$-means clustering to cluster the training queries based on their embeddings. Once the queries are grouped into clusters, we take the centroid embedding of each cluster and input them to Vec2Text to generate one adversarial passage for each cluster. We set $k$ to $10,100,1000$.
Evaluation was then done using NQ test queries. The evaluation measure is $Success@n$ (in line with Zhong et al. (2023)), which represents the percentage of queries for which at least one adversarial passage is retrieved in the top-$n$ results. A higher success rate indicates a greater vulnerability to corpus poisoning.
Table 7 presents corpus poisoning results. When only 10 adversarial passages are generated (grouping training queries with 10-means clustering), both HotFlip and Vec2Text perform poorly on success@10 and 20; however, HotFlip does work when injecting adversarial passages into the top 100 or 1000 rank positions. If the embedding model weights are available, HotFlip is a more potent corpus poisoning attack method when applied to longer rank lists of 100 or 100.
Vec2Text has the advantage of being able to generate many more adversarial passages; when $k=100$ or $k=1000$ HotFlip is too computationally expensive and thus largely impractical. For these cases Vec2Text success rate improves, with $k=1000$ making Vec2Text a considerable threat in corpus poisoning.
The adversarial passages are generated from query embeddings and are such synthetic and may not appear like real passages. Figure 3 shows the two adversarial passages most closely matching a sampled query centroid $\phi_{\mathcal{Q}}$ for both HotFlip and Vec2Text. The cosine similarity with the query centroid embeddings is also reported in the caption. Recall that HotFlip works by flipping tokens one at a time, whereas Vec2Text is a language model decoder. This difference may explain by Vec2Text appears to generate more natural language.
The embeddings of the adversarial passages in Figure 3 still differ from the query centroid embeddings (cosine similarity being 0.96 for HotFlip and 0.85 for Vec2Text). Recall that a perfect Vec2Text would generate an adversarial passage for which the embedding is exactly the same as the query embedding. To understand how this best case scenario affects success@k we run an additional experiment which involved directly inserting the query centroid embeddings into the vector index and then evaluating the success@k. This serves as an upper bound for the HotFlip and Vec2Text corpus poisoning attacks — an analysis overlooked in the original corpus poisoning paper Zhong et al. (2023). Results are shown in Table 7 as "Query centroid (upper bound)". For both different values of $k$ and different success@$n$, the upper bound is much higher than both HotFlip or Vec2Text. This tells us that neither method is optimal yet for the corpus poisoning attack. It also highlights that if a much more effective corpus poising method — one closer to the upper bound — was developed, it could have serious adverse consequences for dense retrievers.
7 Conclusion
Dense retrievers have proven an effective and efficient retrieval method and are now widely adopted in working systems. Much of the benefit comes from using text embeddings to represent and compare information. However, the reliance on text embeddings also opens up dense retrievers to possible threats that exploit such embeddings. Methods like Vec2Text, which can successfully reconstruct the original text from an embedding, could pose serious privacy risks, especially now embeddings are made publicly available via APIs (e.g., OpenAI or Cohere).
This paper reproduces and extends Vec2Text. We begin by uncovering an issue in the original Vec2Text implementation whereby the last hidden layer rather than the dense pooling layer was used for the final text embeddings for one of the considered embedding models. We correct this issue, rerunning the original experiments and showing that Vec2Text is even more effective in retrieval and reconstructibility.
We then apply Vec2text under a number of new conditions: different distance metrics, pooling functions, bottleneck pre-training, embeddings size and model quantization. We discover that the two techniques of mean pooling and bottleneck pre-training increases the risk of reconstructibility. We also show that even zero-shot rankers with poor retrieval effectiveness are vulnerable to Vec2Text reconstruction. A surprising finding was that the compression techniques of embeddings dimension reduction and model product quantization prove to be effective at significantly reducing reconstructibility, while maintaining retrieval effectiveness. Overall, our results give a comprehensive insight into the trade-offs between privacy protection and retrieval effectiveness.
Having clearly outlined the risk, we explore mitigation strategies. We reproduce an existing mitigation strategy that adds noise to the text embeddings, while extending its test to verify if Vec2Text could circumvent this mitigation strategy by training with these noisy embeddings. We show noise injection is a valid mitigation strategy, though comes with trade-offs in terms of retrieval effectiveness and sensitivity to parameter choices. We then propose a new embedding transformation mitigation strategy guaranteed to maintain retrieval effective while completely degrading reconstructibility so privacy is protected. We believe this simple strategy could be easily adopted by services like OpenAI or Cohere to provide their users with an option to make their embeddings more secure.
Finally, we identify that Vec2Text could be applied as a threat to the completely different task for corpus poisoning, whereby adversarial passages are generated and inserted into a corpus such that they are likely be retrieved for any query. Vect2Text poses a great risk here because it can easily generate large numbers of adversarial passages without access to model weights. We show that under certain conditions, corpus poisoning with Vec2Text can pose a serious threat to dense retriever system integrity and user experience.
Overall, this study maps out under what specific conditions Vec2Text could pose a potential threat to current dense retrieval systems, while also highlighting different mitigation strategies and the trade-off between retrieval effectiveness and reconstructibility. While Vec2Text may be a threat, there are some effective and easy to integrate methods to patch current dense retrievers.
Acknowledgments and Disclosure of Funding
We thank Morris et al. for giving permission to re-use their figure illustrating the Vec2Text method. This research is partially funded by Beijing Baidu Netcom Technology Co, Ltd, for the project "Federated Online Learning of Neural Rankers", under funding schema 2022 CCF-Baidu Pinecone.